OpenStudy (anonymous):

Can someone help me out here. I am doing determinants and the question is saying that ABC DEF GHI all equal 4 so given def abc ghi I am supposed to find the determinants!!????

OpenStudy (anonymous):

This is a matrix I just don't know how to put matrix on the computer.

OpenStudy (cruffo):

Do you mean \[det\left(\begin{array}{ccc} A&B&C\\D&E&F\\G&H&I\\ \end{array}\right) = 4\]

OpenStudy (anonymous):

yes

OpenStudy (cruffo):

And you want to know \[det\left(\begin{array}{ccc}D&E&F\\A&B&C\\G&H&I\\\end{array}\right) = ??\]

OpenStudy (anonymous):

yes.

OpenStudy (cruffo):

You multiply the original determinant by -1 for each row exchange.

OpenStudy (anonymous):

so I multiply 4 by -1?

OpenStudy (cruffo):

yep

OpenStudy (anonymous):

so what happens when the next question 2a 2b 2c d e f g h i do i then multiply 4 by 2 and then by -1? Does this make sense?

OpenStudy (cruffo):

Just multiply 4 by 2. You don't have any row exchanges here (from the original matrix), so you don't need to multiply by -1.

OpenStudy (anonymous):

if a row had of been changed i would of had to multiply 4by 2 and then by -1?

OpenStudy (cruffo):

Yep

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

do you have to show your work for it?

OpenStudy (anonymous):

do you have to show your work for it?

OpenStudy (cruffo):

No, it's just using the properties of determinants. If anything you may want to write a sentence explaining why you are doing what you are doing so you don't have to learn it all over again later :)

OpenStudy (cruffo):

If we interchange two rows, the determinant of the new matrix is -1 times the old one. If we multiply one row with a constant, the determinant of the new matrix is the determinant of the old one multiplied by the constant,

OpenStudy (anonymous):

what happens if it is multiplied by two different constants

OpenStudy (cruffo):

you would have the original determinant times the first constant times the second constant.